If the Mica sheet is removed and the plates are moved 1.00 mm farther from each other,

what will be the voltage between the plates, and how much work will it take to move the

plates apart?


1
Expert's answer
2021-02-24T12:47:49-0500

I believe that you are referring to question #166043 in Electric Circuits

and taking the missing values from that question

Initial energy stored"= 252\times10^{-12}J"

Area of the plates = 5cm2

d = "1\times 10^{-3} m"

"c=\dfrac{\epsilon_{\omicron}A}{d}"

"\Rightarrow c=\dfrac{8.85\times10^{-12}\times5\times10^{-4}}{1\times10^{-3}}"

"\Rightarrow c=4.425 pF"


"E=\dfrac{1}{2}cV^2"

"\Rightarrow V^2= \dfrac {252\times10^{-12}\times2}{4.425\times10^{-12}}"

"\Rightarrow V=10.671v"


Work done, "W=QV"

"\Rightarrow W=\dfrac{1}{2}cV^{2}"

"\Rightarrow W=\dfrac{1}{2}\times4.425\times10^{-12}\times(10.671)^{2}"

"\Rightarrow W=2.52\times10^{-12}J"


Energy Stored = Work done